# A Survey of Model Architectures in Information Retrieval

### v2

@@ Abstract
The period from 2019 to the present has represented one of the biggest paradigm shifts in information retrieval (IR) and natural language processing (NLP), culminating in the emergence of powerful large language models (LLMs) from 2022 onward. Methods leveraging pretrained encoder-only models (e.g., BERT) and LLMs have outperformed many previous approaches, particularly excelling in zero-shot scenarios and complex reasoning tasks. This work surveys the evolution of model architectures in IR, focusing on two key aspects: backbone models for feature extraction and end-to-end system architectures for relevance estimation. The review intentionally separates architectural considerations from training methodologies to provide a focused analysis of structural innovations in IR systems. We trace the development from traditional term-based methods to modern neural approaches, particularly highlighting the impact of transformer-based models and subsequent large language models (LLMs). We conclude with a forward-looking discussion of emerging challenges and future directions, including architectural optimizations for performance and scalability, handling of multimodal, multilingual data, and adaptation to novel application domains such as autonomous search agents that is beyond traditional search paradigms.
@@ 1 Introduction
Figure 1: An overview of this survey.
@@
Information Retrieval (IR) aims to retrieve relevant information sources to satisfy users’ information needs. In the past decades, IR has become indispensable for efficiently and effectively accessing vast amounts of information across various applications. Beyond its traditional role, IR now also plays a critical role in assisting large language models (LLMs) to generate grounded and factual responses. Research in IR primarily centers on two key aspects: (1) extracting better query and document feature representations, and (2) developing more accurate relevance estimators. The approaches for extracting query and document features have evolved from traditional term-based methods, such as boolean logic (Radecki, 1979; Kraft & Buell, 1983) and vector space models (Salton et al., 1975), to modern solutions such as dense retrieval based on pre-trained language models (Lee et al., 2019; Karpukhin et al., 2020; Logeswaran et al., 2019; Lin et al., 2022, inter alia).
@@
Relevance estimators have evolved alongside advances in feature representations. Early approaches, including probabilistic and statistical language models, computed relevance with simple similarity functions based on term-based features. Learning-to-rank (LTR) techniques later emerged, incorporating machine learning models like support vector machines (Cortes & Vapnik, 1995), boosting methods (Kearns & Valiant, 1994; Freund & Schapire, 1995) as well as multi-layer neural networks for relevance estimation (Li, 2011). The success of LTR methods can be largely attributed to their extensive use of manually engineered features, derived from both statistical properties of text terms and user behavior data collected from web browsing traffic (Qin & Liu, 2013). In the 2010s, a vast literature explored neural rerankers in different architectures to capture the semantic similarity between queries and documents (Pang et al., 2016; Guo et al., 2016a; Xiong et al., 2017; Dai et al., 2018, inter alia). Then pre-trained transformers, represented by BERT (Devlin et al., 2019) and its variants (Liu, 2019; Sun et al., 2019; Lan et al., 2020; Beltagy et al., 2020), quickly revolutionized the model design, leading to an era where retrieval and ranking models adopt simpler architectures for relevance estimation, such as dot product operations and MLP prediction heads, which operate on learned neural representations (MacAvaney et al., 2019; Lee et al., 2019; Karpukhin et al., 2020; Nogueira et al., 2020; Lin et al., 2022; Formal et al., 2021b; a).
@@
Recent advancements in LLMs have revolutionized applied machine learning (ML) communities, including IR. One intriguing property of modern instruction-following LLMs (e.g., ChatGPT OpenAI (2022)) is that they can be used for feature extraction and relevance estimation, achieving strong performance without extensive training (Ni et al., 2022a; Neelakantan et al., 2022; BehnamGhader et al., 2024; Sun et al., 2023; Qin et al., 2024b, inter alia). The rise of these models builds upon a rich foundation of neural architectures, including the classical Transformer architecture with multi-head attention (MHA, Vaswani et al., 2017), Recurrent Neural Networks (RNN, Elman, 1990), Attention mechanisms (Bahdanau, 2014), and pre-trained static word representations like Word2Vec (Mikolov, 2013) and GloVe (Pennington et al., 2014), as well as (Collobert et al., 2011; Le & Mikolov, 2014, inter alia).
@@
This work reviews the evolution of model architectures in IR (with an overview in Fig.˜1). Here, the meaning of model architecture is twofold: it describes (1) backbone models for extracting query and document feature representations, and (2) end-to-end system architectures that process raw inputs, perform feature extraction, and estimate relevance. Different from prior works and surveys (Lin et al., 2022; Zhu et al., 2023), we intentionally separate our discussion of model architectures from training methodologies and deployment best practices to provide a focused architectural analysis. The shift towards neural architectures, particularly Transformer-based models, has fundamentally transformed IR by enabling rich, contextualized representations and improved handling of complex queries. While this evolution has enhanced retrieval precision, it also presents new challenges, especially with the emergence of LLMs. These challenges include the need for architectural innovations to optimize performance and scalability, handle multimodal and multilingual data, and understand complex instructions. Moreover, as IR systems are increasingly integrated into diverse applications — from robotics (Xie et al., 2024), protein structure discovery (Jumper et al., 2021) to autonomous agents (Wu et al., 2023a; Chen et al., 2025; Hu et al., 2025; OpenAI, 2025; Wu et al., 2025b, inter alia) that are capable of reasoning and search — the field must evolve beyond traditional search paradigms. We conclude this survey by examining these challenges and discussing their implications for the future of IR model architectures research.
@@ 2 Background and Terminology
Given a query $\mathcal{Q}$, the task is to find a ranked list of $k$ documents, denoted as $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, that exhibit the highest relevance to $\mathcal{Q}$. This is achieved either by retrieving top-$k$ documents from a large collection $\mathcal{C}$ ($|\mathcal{C}|\gg|k|$ ), which typically comprises millions or billions of documents, or by reranking the top-$k$ candidates returned by a retriever. System performance is measured using standard, list-wise IR metrics. Common metrics include:
@@
The Multi-Stage "Retrieve-then-Rerank" Architecture.
Modern large-scale IR systems almost universally operate on a multi-stage pipeline, commonly known as the "retrieve-then-rerank" architecture. This design balances the trade-off between efficiency and effectiveness.
@@
    Retrieval (or First-Stage Ranking): In the first stage, a computationally efficient but less precise model scans the entire collection $\mathcal{C}$ (potentially billions of documents) to quickly identify an initial set of several hundred or thousand candidate documents. These models, often called retrievers, must be extremely fast. Examples include traditional models like BM25 (§˜3) or modern bi-encoder models (§˜6).
@@
    Reranking (or Second-Stage Ranking): In the second stage, a more powerful but computationally expensive model, known as a reranker, is applied only to the small candidate set returned by the retriever. This model can afford to perform deep, fine-grained analysis of the interaction between the query and each candidate document to produce a more accurate final ranking. Examples include Learning-to-Rank models (§˜4) and modern cross-encoder transformers (§˜6).
@@
A query expresses an information need and serves as input to the ad hoc retrieval system. We denote document as the atomic unit for retrieval and ranking. Our discussions are primarily based on text-based documents, but a documnent can also refer to a webpage or an email, depending on the actual IR application of interest.
@@ 3 Traditional IR Models
In this section, we briefly review traditional Information Retrieval (IR) models prior to neural methods, with a focus on the Boolean model, vector space model, probabilistic model, and statistical language model. These models, which serve as the foundation for later developments in IR (§§˜4, 5, 6 and 7), are built upon the basic unit of a “term” in their representations (Nie, 2010).
@@
In Vector Space Models (VSMs, Salton et al., 1975), the queries and documents are represented by vectors, e.g., $\mathcal{Q}=<q_{1},q_{2},\dots,q_{n}>$ and $\mathcal{D}=<d_{1},d_{2},\dots,d_{n}>$. The vector space is defined by a vocabulary of terms $\mathcal{V}=<t_{1},t_{2},\dots,t_{n}>$ and each element ($q_{i}$ or $d_{i}$, $1\leq i\leq n$) in the vectors represents the weight of the corresponding term in the query or the document. The weights $q_{i}$ or $d_{i}$ could be binary, representing presence or absence. Given the vector representations, the relevance score is estimated by a similarity function between the query $\mathcal{Q}$ and the document $\mathcal{D}$. The weights $q_{i}$ or $d_{i}$ can be determined by more sophisticated schema (Salton & Buckley, 1988), such as TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 1995). This allows for more abundant features that can improve the capacity and accuracy of the models. Besides, given the vector representations of query $\mathcal{Q}$ and document $\mathcal{D}$, the most commonly used is cosine similarity, defined as:
@@
The smoothing mechanisms (Baeza-Yates et al., 1999) are necessary to deal with zero occurrences of $t_{i}$. Except for the binary independence retrieval model, more sophisticated probabilistic models have been proposed in the literature (Wong & Yao, 1989; Fuhr, 1992), such as the inter-dependency between terms (Van Rijsbergen, 1979).
@@ 4 Learning-to-Rank Model Architectures
Different from traditional IR models that rely on heuristic-based scoring formulas (§˜3), Learning-to-Rank (LTR) reframes ranking as a supervised machine learning problem (Liu, 2009). The core idea is to train a model that can optimally combine a wide array of signals, or "features," to predict the relevance of documents to a query. For each query-document pair $(\mathcal{Q}_{i},\mathcal{D}_{i})$, a $k$-dimensional feature vector $\mathbf{x}_{i}\in\mathbb{R}^{k}$ is extracted, and a relevance label $\mathbf{y}_{i}$ (e.g., from human judgments) is provided. The goal is to learn a ranking model $f$ parameterized by $\theta$ that minimizes an empirical loss $l(\cdot)$ on a labeled training set $\Psi$:
@@
Name Model Backbone Architecture Loss Function
MART (Friedman, 2001) ML Boosting Pointwise
RankBoost (Freund et al., 2003) ML Boosting Pairwise
RankNet (Burges et al., 2005) Neural Nets Neural Network Pairwise
RankSVM (Joachims, 2006) ML SVM Pairwise
LambdaRank (Burges et al., 2006) Neural Nets Neural Network Pairwise
ListNet (Cao et al., 2007) Neural Nets Neural Network Listwise
SoftRank (Taylor et al., 2008) Neural Nets Neural Network Listwise
ListMLE (Xia et al., 2008) ML Linear Listwise
LambdaMART (Burges, 2010) ML GBDT Listwise
ApproxNDCG (Qin et al., 2010) ML Linear Listwise
DLCM (Ai et al., 2018a) Neural Nets Neural Network Listwise
GSF (Ai et al., 2019) Neural Nets Neural Network Listwise
ApproxNDCG (Bruch et al., 2019) Neural Nets Neural Network Listwise
SetRank (Pang et al., 2020) Neural Nets Self Attention Blocks Listwise
@@
Pointwise, Pairwise, and Listwise Approaches with ML Models.
@@
The pairwise approach addresses this by focusing on the relative order of document pairs. Given two documents $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ for the same query, the goal is to predict which one is more relevant. This transforms ranking into a binary classification problem. Seminal pairwise models include RankSVM (Joachims, 2006), which adapts the Support Vector Machine framework to maximize the number of correctly ordered pairs, and RankNet (Burges et al., 2005), which uses a neural network and a probabilistic cost function based on pairwise logistic loss. While more aligned with the nature of ranking than the pointwise approach, pairwise methods still do not directly optimize the list-based evaluation metrics (e.g., NDCG, ERR) that are standard in IR evaluation.
The listwise approach directly tackles this issue by defining the loss function over an entire list of documents for a given query. These methods aim to directly optimize ranking metrics. A pivotal line of work began with LambdaRank (Burges et al., 2006), which observed that for gradient-based optimization, one only needs the gradients of the loss function. It introduced "Lambda gradients," which are derived from the change in an IR metric (like NDCG) when two documents in a ranked list are swapped. This technique was then combined with Multiple Additive Regression Trees (MART), a Gradient Boosted Decision Tree (GBDT) algorithm (Friedman, 2001), to create LambdaMART (Wu et al., 2010). Due to its strong performance and robustness, LambdaMART became a dominant industry standard for many years (Ke et al., 2017).
Neural LTR Models and Other Directions.

### v3

@@ Abstract
The period from 2019 to the present marks one of the most significant paradigm shifts in information retrieval (IR) and natural language processing (NLP), culminating in the emergence of powerful large language models (LLMs) from 2022 onward. Methods based on pretrained encoder-only architectures (e.g., BERT) as well as decoder-only generative LLMs have outperformed many earlier approaches, demonstrating particularly strong performance in zero-shot scenarios and complex reasoning tasks. This survey examines the evolution of model architectures in IR, with a focus on two key aspects: backbone models for feature extraction and end-to-end system architectures for relevance estimation. To maintain analytical clarity, we deliberately separate architectural design from training methodologies, enabling a focused examination of structural innovations in IR systems. We trace the progression from traditional term-based retrieval models to modern neural approaches, highlighting the transformative impact of transformer-based architectures and subsequent LLM developments. The survey concludes with a forward-looking discussion of open challenges and emerging research directions, including architectural optimization for efficiency and scalability, robust handling of multimodal and multilingual data, and adaptation to novel application domains such as autonomous search agents, which may represent the next paradigm in IR.
@@ 1 Introduction
Figure 1: High-level overview of the survey structure and scope.
@@
Information Retrieval (IR) aims to retrieve relevant information sources to satisfy users’ information needs. In the past decades, IR has become indispensable for efficiently and effectively accessing vast amounts of information across various applications. Beyond its traditional role, IR now also plays a critical role in assisting large language models (LLMs) to generate grounded and factual responses under the generative AI era. Research in IR primarily centers on two key aspects: (1) extracting better query and document feature representations, and (2) developing more accurate relevance estimators. Extracting better query and document feature representations focuses on modeling textual content so that queries and documents can be compared in a meaningful space, ranging from early term-frequency vectors (e.g., TF–IDF (Sparck Jones, 1972) and BM25 (Robertson & Zaragoza, 2009)) to modern contextual embeddings derived from pre-trained language models. Developing more accurate relevance estimators then builds on these representations to assess how well a document satisfies an information need, using scoring functions or learned ranking models such as BM25, neural interaction models, or later learning-to-rank frameworks that combine multiple signals. The approaches for extracting query and document features have evolved from traditional term-based methods, such as boolean logic (Radecki, 1979; Kraft & Buell, 1983) and vector space models (Salton et al., 1975), to modern solutions such as dense retrieval based on pre-trained language models (Lee et al., 2019; Karpukhin et al., 2020; Logeswaran et al., 2019; Lin et al., 2022, inter alia).
@@
Relevance estimators have evolved alongside advances in feature representations. Early approaches, including probabilistic and statistical language models, computed relevance with simple similarity functions based on term-based features. Learning-to-rank (LTR) techniques later emerged, incorporating machine learning models like support vector machines (Cortes & Vapnik, 1995), boosting methods (Kearns & Valiant, 1994; Freund & Schapire, 1995) as well as multi-layer neural networks for relevance estimation (Li, 2011). The success of LTR methods can be largely attributed to their extensive use of manually engineered features, derived from both statistical properties of text terms and user behavior data collected from web browsing traffic (Qin & Liu, 2013). In the 2010s, a vast literature explored neural rerankers in different architectures to capture the semantic similarity between queries and documents (Pang et al., 2016; Guo et al., 2016a; Xiong et al., 2017; Dai et al., 2018, inter alia). Then pre-trained transformers, represented by BERT (Devlin et al., 2019) and its variants (Liu et al., 2019; Sun et al., 2019; Lan et al., 2020; Beltagy et al., 2020), quickly revolutionized the model design, leading to an era where retrieval and ranking models adopt simpler architectures for relevance estimation, such as dot product operations and MLP prediction heads, which operate on learned neural representations (MacAvaney et al., 2019b; Lee et al., 2019; Karpukhin et al., 2020; Nogueira et al., 2020; Lin et al., 2022; Formal et al., 2021b; a).
@@
Recent advancements in LLMs have revolutionized applied machine learning (ML) communities, including IR. One intriguing property of modern instruction-following LLMs (e.g., ChatGPT OpenAI (2022)) is that they can be used for feature extraction and relevance estimation, achieving strong performance without extensive training (Ni et al., 2022a; Neelakantan et al., 2022; BehnamGhader et al., 2024; Sun et al., 2023; Qin et al., 2024b, inter alia). The rise of these models builds upon a rich foundation of neural architectures, including the classical Transformer architecture with multi-head attention (MHA, Vaswani et al., 2017), Recurrent Neural Networks (RNN, Elman, 1990), attention mechanisms (Bahdanau et al., 2014), and pre-trained static word representations like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) (Collobert et al., 2011; Le & Mikolov, 2014, inter alia).
@@
This work reviews the evolution of model architectures in IR (with an overview in Figure˜1). Here, the meaning of model architecture is twofold: it describes (1) backbone models for extracting query and document feature representations, and (2) end-to-end system architectures that process raw inputs, perform feature extraction, and estimate relevance. Different from prior works and surveys (Lin et al., 2022; Zhu et al., 2023), we intentionally separate our discussion of model architectures from training methodologies and deployment best practices to deliver a focused architectural analysis. This analysis centers on the core components of AI infrastructure in the LLM era. The shift towards neural architectures, particularly Transformer-based models, has fundamentally transformed IR by enabling rich, contextualized representations and improved capacity for handling complex queries. While this evolution enhanced retrieval performance, it also presents new challenges, especially with the development of LLMs. These challenges include the need for architectural innovations to optimize performance and scalability, handle multimodal and multilingual data, and understand complex instructions. Moreover, as IR systems are increasingly integrated into diverse applications — from robotics (Xie et al., 2024b), protein structure discovery (Jumper et al., 2021) to autonomous agents (Wu et al., 2023a; Chen et al., 2025; Hu et al., 2025; OpenAI, 2025; Wu et al., 2025b, inter alia) that are capable of reasoning and search — the field must evolve beyond traditional search paradigms. We conclude this survey by examining these challenges and discussing their implications for the future of IR model architectures research.
@@ 2 Background and Terminology
Given a query $\mathcal{Q}$, the task is to find a ranked list of $k$ documents, denoted as $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, that exhibit the highest relevance to $\mathcal{Q}$. This is achieved either by retrieving top-$k$ documents from a large collection $\mathcal{C}$ ($|\mathcal{C}|\gg k$), which typically comprises millions or billions of documents, or by reranking the top-$k$ candidates returned by a retriever. System performance is measured using standard, list-wise IR metrics. Common metrics include:
@@
The Multi-Stage “Retrieve-then-Rerank” Architecture.
Modern large-scale IR systems almost universally operate on a multi-stage pipeline, commonly known as the “retrieve-then-rerank” architecture. This design balances the tradeoff between efficiency and effectiveness.
@@
    Retrieval (or First-Stage Ranking): In the first stage, a computationally efficient but less precise model scans the entire collection $\mathcal{C}$ (potentially billions of documents) to quickly identify an initial set of several hundred or thousand candidate documents. These models, often called retrievers, must be extremely fast. Examples include traditional models like BM25 (Section˜3) or modern bi-encoder models (Section˜6).
@@
    Reranking (or Second-Stage Ranking): In the second stage, a more powerful but computationally expensive model, known as a reranker, is applied only to the small candidate set returned by the retriever. This model can afford to perform deep, fine-grained analysis of the interaction between the query and each candidate document to produce a more accurate final ranking. Examples include Learning-to-Rank models (Section˜4) and modern cross-encoder transformers (Section˜6).
@@
A query expresses an information need and serves as input to the ad hoc retrieval system. We denote document as the atomic unit for retrieval and ranking. Our discussions are primarily based on text-based documents, but a document can also refer to a webpage or an email, depending on the actual IR application of interest.
@@ 3 Traditional IR Models
In this section, we briefly review traditional Information Retrieval (IR) models prior to neural methods, with a focus on the Boolean model, vector space model, probabilistic model, and statistical language model. These models, which serve as the foundation for later developments in IR (Sections˜4, 5, 6 and 7), are built upon the basic unit of a “term” in their representations (Nie, 2010).
@@
In Vector Space Models (VSMs, Salton et al., 1975), the queries and documents are represented by vectors, e.g., $\mathcal{Q}=<q_{1},q_{2},\dots,q_{n}>$ and $\mathcal{D}=<d_{1},d_{2},\dots,d_{n}>$. The vector space is defined by a vocabulary of terms $\mathcal{V}=<t_{1},t_{2},\dots,t_{n}>$ and each element ($q_{i}$ or $d_{i}$, $1\leq i\leq n$) in the vectors represents the weight of the corresponding term in the query or the document. The weights $q_{i}$ or $d_{i}$ could be binary, representing presence or absence. Given the vector representations, the relevance score is estimated by a similarity function between the query $\mathcal{Q}$ and the document $\mathcal{D}$. The weights $q_{i}$ or $d_{i}$ can be determined by more sophisticated schema (Salton & Buckley, 1988), such as TF-IDF (Sparck Jones, 1972) and BM25 (Robertson & Zaragoza, 2009; Robertson et al., 1995). This allows for more abundant features that can improve the capacity and accuracy of the models. Besides, given the vector representations of query $\mathcal{Q}$ and document $\mathcal{D}$, the most commonly used is cosine similarity, defined as:
@@
The smoothing mechanisms (Baeza-Yates & Ribeiro-Neto, 1999) are necessary to deal with zero occurrences of $t_{i}$. Except for the binary independence retrieval model, more sophisticated probabilistic models have been proposed in the literature (Wong & Yao, 1989; Fuhr, 1992), such as the inter-dependency between terms (Van Rijsbergen, 1979).
@@ 4 Learning-to-Rank Model Architectures
Different from traditional IR models that rely on heuristic-based scoring formulas (Section˜3), Learning-to-Rank (LTR) frames ranking as a supervised machine learning problem (Liu, 2009). The core idea is to train a model that can optimally combine a wide array of signals, or “features”, to predict the relevance of documents to a query. For each query-document pair $(\mathcal{Q}_{i},\mathcal{D}_{i})$, a $k$-dimensional feature vector $\mathbf{x}_{i}\in\mathbb{R}^{k}$ is extracted, and a relevance label $\mathbf{y}_{i}$ (e.g., from human judgments) is provided. The goal is to learn a ranking model $f$ parameterized by $\theta$ that minimizes an empirical loss $l(\cdot)$ on a labeled training set $\Psi$:
@@
Name Backbone Architecture Loss Function
MART (Friedman, 2001) Boosting Pointwise
RankBoost (Freund et al., 2003) Boosting Pairwise
RankNet (Burges et al., 2005) Neural Network Pairwise
RankSVM (Joachims, 2006) SVM Pairwise
LambdaRank (Burges et al., 2006) Neural Network Pairwise
ListNet (Cao et al., 2007) Neural Network Listwise
SoftRank (Taylor et al., 2008) Neural Network Listwise
ListMLE (Xia et al., 2008) Linear Listwise
LambdaMART (Burges, 2010) GBDT Listwise
ApproxNDCG (Qin et al., 2010) Linear Listwise
DLCM (Ai et al., 2018a) Neural Network Listwise
GSF (Ai et al., 2019) Neural Network Listwise
ApproxNDCG (Bruch et al., 2019) Neural Network Listwise
SetRank (Pang et al., 2020) Self Attention Blocks Listwise
4.1 Pointwise, Pairwise, and Listwise Approaches with ML Models
@@
The pairwise approach addresses this by focusing on the relative order of document pairs. Given two documents $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ for the same query, the goal is to predict which one is more relevant. This transforms ranking into a binary classification problem. Seminal pairwise models include RankSVM (Joachims, 2006), which adapts the Support Vector Machine framework to maximize the number of correctly ordered pairs, and RankNet (Burges et al., 2005), which uses a neural network and a probabilistic cost function based on pairwise logistic loss. While more aligned with the nature of ranking than the pointwise approach, pairwise methods still do not directly optimize the list-based evaluation metrics (e.g., nDCG, MRR) that are standard in IR evaluation.
The listwise approach directly tackles this issue by defining the loss function over an entire list of documents for a given query. These methods aim to directly optimize ranking metrics. A pivotal line of work began with LambdaRank (Burges et al., 2006), which observed that for gradient-based optimization, one only needs the gradients of the loss function. It introduced “Lambda gradients”, which are derived from the change in an IR metric (like nDCG) when two documents in a ranked list are swapped. This technique was then combined with Multiple Additive Regression Trees (MART), a Gradient Boosted Decision Tree (GBDT) algorithm (Friedman, 2001), to create LambdaMART (Wu et al., 2010). Due to its strong performance and robustness, LambdaMART became a dominant industry standard for many years (Ke et al., 2017).
4.2 Neural LTR Models